OpenStudy (snivvles.com):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

Nnesha (nnesha):

let sinx = y \[\huge\rm 7y^2-14y+2=-5\] now you have quadratic equation set it equal to zero (move -5 to the left side)

OpenStudy (snivvles.com):

so \[7y ^{2}-14y+7=0\]?

Nnesha (nnesha):

yes right now can you solve for y ?

OpenStudy (snivvles.com):

\[\frac{ 14+-\sqrt{196-4(7)(7)} }{ 2(7) }\]

Nnesha (nnesha):

oh wait you can take out the common factor

Nnesha (nnesha):

\[\huge\rm 7y ^{2}-14y+7=0\] what is gcf ?

OpenStudy (snivvles.com):

7 so \[\frac{ 2+-\sqrt{4-4} }{ 2 }\]

Nnesha (nnesha):

right!

OpenStudy (snivvles.com):

so \[\frac{ 2 }{ 2 }\]

OpenStudy (snivvles.com):

then I find where that happens on the unit circle, right?

OpenStudy (snivvles.com):

but it simplifies to \[\frac{ 1 }{ 1 }\]

Nnesha (nnesha):

yes right so \[\large\rm \frac{ 2 +0 }{ 2} ~and~ \frac{ 2-0 }{ 2 }\] both equal to one y= 1 now you can replace y with sinx sin x =1 sin x =1

OpenStudy (snivvles.com):

so \[\frac{ \pi }{ 2 }\]

OpenStudy (snivvles.com):

right?

Nnesha (nnesha):

find `all solution`

Nnesha (nnesha):

wait do you need in radians or degree ?

OpenStudy (snivvles.com):

so 3pi/2 5pi/2 7pi/2 etc. ?

OpenStudy (snivvles.com):

radians

Nnesha (nnesha):

yes pi /2 is right !

OpenStudy (snivvles.com):

is 3pi/2 right?

Nnesha (nnesha):

sinx = positive one

OpenStudy (snivvles.com):

oh right so is 5pi/2 right?

OpenStudy (snivvles.com):

so like pi/2 5pi/2 9pi/2 etc.

OpenStudy (snivvles.com):

right?

Nnesha (nnesha):

yes 5pi/2 can be the solutions

OpenStudy (snivvles.com):

Ok, thanks!

Nnesha (nnesha):

but remember sinx =1 (multiplicity of 2)(sinx= 1 , sinx =1)

Nnesha (nnesha):

np good job!

OpenStudy (snivvles.com):

OK!